# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

### v1

@@ 4.1 NFDM with optimal transport
DDPM (${\mathcal{L}_{\mathrm{simple}}}$) (Ho et al., 2020) $1000$ $3.17$
DDPM (ELBO) (Ho et al., 2020) $1000$ $13.51$
Flow Matching (Lipman et al., 2022) $142$ $6.35$ $122$ $5.02$ $138$ $14.14$
DDIM (Song et al., 2020a) $10$ $13.36$
@@
DPM Solver (Lu et al., 2022) $24$ $2.75$
Trajectory Curvature Minimization (Lee et al., 2023) $5$ $18.74$
@@
Multisample Flow Matching (Pooladian et al., 2023) $12$ $7.18$ $12$ $17.6$
@@
5 Experiments
We first showcase results demonstrating that NFDM consistently achieves better likelihood compared to baselines, obtaining state-of-the-art diffusion modeling results on the CIFAR-10 and downsampled ImageNet datasets. Then, we explore the NFDM-OT modification, which penalizes the curvature of the deterministic generative trajectories. The NFDM-OT reduces trajectory curvature, significantly lowering the number of generative steps required for sampling.
We evaluate NFDM on synthetic data, CIFAR-10 (Krizhevsky et al., 2009), and two downsampled ImageNet (Deng et al., 2009; Van Den Oord et al., 2016) datasets. For parameterization we use multi-layer perceptrons for synthetic data and the U-Net architecture from Dhariwal & Nichol (2021) for images.
We report the NLL in Bits Per Dimension (BPD) and sample quality measured by the FID score (Heusel et al., 2017). The NLL is calculated by integrating the ODEs using the RK45 solver Dormand & Prince (1980), with all NLL metrics computed on test data. For the FID, we provide the average over 50k images.
For a detailed description of parameterizations and other experimental details, please refer to Appendix C.
(a) Score SDE (Song et al., 2020b)
(b) NFDM (this paper)
Figure 1: Comparison of trajectories between the data distribution (on the left) and the prior distribution (on the right).
5.1 Likelihood estimation
We have trained NFDM on the CIFAR-10 and downsampled ImageNet datasets. The NLL results are summarized in Table 1. Notably, NFDM outperforms diffusion-based baselines on all three datasets. We consider this enhanced performance of the NFDM a natural progression due to the following reasons.
First, it is well-established (Song et al., 2021; Zheng et al., 2023b) that diffusion models exhibit improved likelihood estimation when trained with the full ELBO objective. As delineated in Section 3.3, the objective $\mathcal{L}$ (eq. 18) used for training the NFDM constitutes a variational bound on the likelihood.
Second, as previously mentioned, diffusion models can be seen as hierarchical VAEs. From this perspective, the baselines in Table 1 resemble VAEs with either fixed or constrained variational distributions. In contrast, the NFDM extends beyond these baselines by providing a more flexible variational distribution. This flexibility allows the NFDM to better conform to the reverse process, consequently enhancing likelihood estimation.
5.2 Straight trajectories
We next evaluate NFDM-OT, which is designed to penalize the curvature of deterministic generative trajectories. Initially, we compare NFDM-OT with a conventional continuous-time diffusion model. Figure 1(a) illustrates deterministic trajectories between a two-dimensional data distribution and a unit Gaussian distribution learnt by a conventional diffusion model (Song et al., 2020b). Figure 1(b) depicts trajectories learnt by NFDM-OT. Conventional diffusion, being constrained in its forward process, learns highly curved trajectories, whereas NFDM-OT successfully learns straight generative trajectories as desired.
Subsequently, we present evaluation results of NFDM-OT on image datasets. Table 2 reports the FID scores for $2$, $4$, and $12$ Number of Function Evaluations (NFE) with respect to the function $\hat{f}_{\theta,\varphi}$ (eq. 16). In this experiment, we employ Euler’s method for sampling integration. For the specified NFEs, NFDM-OT demonstrates superior sample quality compared to other approaches with similar NFE values. Specifically, NFDM-OT outperforms approaches that minimize the curvature of generative trajectories (Pooladian et al., 2023; Lee et al., 2023).
Importantly, NFDM-OT is trained with an ELBO-based objective (eq. 18), which is known to yield higher FID scores for diffusion models (Song et al., 2021; Zheng et al., 2023b). In contrast, some of the approaches listed in Table 2 are trained with different objectives, leading to improved FID scores. Even so, for comparable NFE values NFDM-OT still achieves superior results.
The primary aim of the experiments with NFDM-OT is not to introduce a novel model that surpasses others in few-step generation, but rather to showcase NFDM’s ability to learn generative dynamics with specific properties. The straightness of the trajectories is just one example of such properties. We leave it for future research to explore different parameterizations and a modified objectives for NFDM that may yield even better results.
@@ 6 Related work
Diffusion models, originally proposed by Sohl-Dickstein et al. (2015), have evolved significantly through subsequent developments (Song & Ermon, 2019; Ho et al., 2020). These advancements have resulted in remarkable generative quality for high-dimensional data distributions (Dhariwal & Nichol, 2021; Saharia et al., 2022). Nevertheless, conventional diffusion models, typically relying on a linear Gaussian forward process, may not optimally fit some data distributions. To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection (Rissanen et al., 2022; Daras et al., 2022; Hoogeboom & Salimans, 2022), diffusion in the wavelet spectrum (Phung et al., 2023), and forward processes based on the exponential family (Okhotin et al., 2023). These models are limited by their fixed forward processes and may be seen as specific instances of NFDM.
@@
Some studies have focused on making the forward process learnable. Approaches include a learnable noise injection schedule (Kingma et al., 2021; Nichol & Dhariwal, 2021; Sahoo et al., 2023) and learning data transformations like time-independent transformations based on VAEs (Vahdat et al., 2021; Rombach et al., 2022) and normalizing flows (Kim et al., 2022), or time-dependent transformations (Gu et al., 2022; Nielsen et al., 2023; Bartosh et al., 2023). These methods can also be considered special cases of NFDM with specific transformations $F_{\varphi}$ (eq. 9).
Recent studies (Zhang & Chen, 2021; De Bortoli et al., 2021; Shi et al., 2023) have explored generative models based on Schrödinger Bridge theory and finite-time diffusion constructions. While these models offer novel approaches to learning forward transformations, in contrast to NFDM, they depart from the simulation-free paradigm by requiring the full simulation of the stochastic processes for training, increasing their computational cost.
@@
Despite the simulation-free nature of the training procedure, diffusion models still necessitate full reverse process simulations for sample generation, leading to slow and computationally expensive inference. To address this, in an orthogonal line of works alternative sampling methods have been studied, such as deterministic sampling in discrete (Song et al., 2020a) and continuous time (Song et al., 2020b) and novel numerical solvers (Tachibana et al., 2021; Liu et al., 2022a; Lu et al., 2022; Shaul et al., 2023b). Additionally, distillation techniques have been applied to both discrete (Salimans & Ho, 2022) and continuous time models (Song et al., 2023; Zheng et al., 2023a; Yin et al., 2023) to enhance sampling speed, albeit at the cost of training additional models. Most of these approches are compatible with NFDM and may be combined for further gains.
The significance of deterministic trajectories in efficient sampling was highlighted by Karras et al. (2022). At the same time, building on diffusion model concepts, Lipman et al. (2022); Liu et al. (2022b); Albergo & Vanden-Eijnden (2022) introduced simulation-free methods for learning deterministic generative dynamics. Based on these ideas, Liu et al. (2022b) and Liu (2022) proposed distillation procedures to straighten deterministic generative trajectories. Since these methods are based on distillation, we consider them orthogonal to NFDM.
Pooladian et al. (2023) and Tong et al. (2023) proposed to construct the forward process with optimal data-noise couplings to learn straighter generative trajectories. While this method is theoretically justified, it relies on estimating an optimal coupling over minibatches of the entire dataset, which, for large datasets, may become uninformative as to the true coupling. In contrast, NFDM-OT directly penalizes curvature of generative trajectories.
Lee et al. (2023) suggested learning the forward process distribution $q(\mathbf{z}_{1}|\mathbf{x})$ and interpolating linearly between $\mathbf{x}$ and $\mathbf{z}_{1}$. Similarly, Shaul et al. (2023a) aimed to modify the forward process to reduce curvature in the generative process. Both approaches can be viewed as specific instances of NFDM.
Lastly, Albergo et al. (2023) suggested optimizing the forward process simultaneously to training the reverse process, resultin in a challenging high dimentional min-max problem. Moreover, only the theoretical algorithm was presented, without empirical study and descriptions of ways to parameterize such a learnable forward process.
We continue discussing related works in Appendix B.
7 Discussion and Limitations
@@
In this paper we introduced NFDM, a novel simulation-free framework for improved diffusion modeling through a learnable forward processes. Our approach outperforms baseline diffusion models on standard benchmarks, showcasing the effectiveness of NFDM in log-likelihood estimation.
We introduce NFDM not merely as a specific model, but rather as a versatile framework that facilitates the predefined specification and learning of the forward process. For the sake of simplicity in this work, we have chosen a straightforward Gaussian parameterization for the forward process (see Section 3.5). However, NFDM is capable of supporting far more flexible parameterizations beyond standard Gaussian distributions.
Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 4.1 represents just one specific example of the restrictions that NFDM can accommodate. NFDM opens up significant prospects for exploring new generative dynamics.
@@
Nonetheless, the advantages of NFDM come with certain trade-offs. Once the forward process is parameterized using a neural network, this leads to increased computational costs compared to conventional diffusion models. In our experiments, an optimization iteration of NFDM takes approximately 2.2 times longer than that of conventional diffusion models.
@@
Despite these challenges, we are optimistic about the potential of NFDM in various domains and practical applications, given its flexibility in learning generative processes. We also believe that with alternative parameterizations, modifications of the objective, and the integration of orthogonal approaches like distillation, NFDM has the potential to achieve even better results.
@@ Appendix A Derivations
A.1 Derivation of objective
For the forward process, we have the conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which is implicitly defined by the invertible transformation $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x})$ (eq. 9), where $\varepsilon\sim q(\varepsilon)$. Additionally, there’s a SDE (eq. 15) starting from $\mathbf{z}_{1}\sim q(\mathbf{z}_{1}|\mathbf{x})$ and flowing backwards in time:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\mathbf{w}.$ (27)
This SDE corresponds to the marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$.
The reverse process involves a prior distribution $p(\mathbf{z}_{1})$ and a reverse SDE (eq. 15):
@@ A.1 Derivation of objective
Our goal is to derive a variational bound on the negative log-likelihood of the model.
We begin by discretizing the conditional reverse SDE in eq. 27 and the reverse SDE in eq. 28, transitioning from continuous-time trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ to discrete-time trajectories $\bar{\mathbf{z}}_{0},\bar{\mathbf{z}}_{\frac{1}{T}},\dots,\bar{\mathbf{z}}_{1}$, where $T$ represents the number of discrete steps. For convenience, we also introduce $\Delta t=\frac{1}{T}$.
The discretization of the conditional reverse SDE is as follows:
$\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})=q_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x}),\quad\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (29)
We similarly discretize the reverse SDE:
$\bar{p}(\bar{\mathbf{z}}_{1})=p(\bar{\mathbf{z}}_{1}),\quad\bar{p}(\mathbf{x}|\bar{\mathbf{z}}_{0})=p(\mathbf{x}|\bar{\mathbf{z}}_{0}),\quad\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (30)
@@
Now, considering these discrete processes and recognizing that they are Markovian conditional and unconditional processes moving backwards in time, we can leverage a key result from Ho et al. (2020):
$\mathbb{E}_{q(\mathbf{x})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x})\right]\leq\bar{\mathcal{L}}_{\mathrm{recon}}+\bar{\mathcal{L}}_{\mathrm{diff}}+\bar{\mathcal{L}}_{\mathrm{prior}},\quad\textrm{where}$ (31)
¯$\bar{\mathcal{L}}_{\mathrm{recon}}=\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{0})}\left[-\log p(\mathbf{x}|\bar{\mathbf{z}}_{0})\right],$ (32)
@@
Consequently, eq. 31 establishes a variational upper bound on the negative log-likelihood of discretized reverse process. Taking the limits of this upper bound, the reconstruction term ¯$\bar{\mathcal{L}}\mathrm{recon}$ and the prior term ¯$\bar{\mathcal{L}}\mathrm{prior}$ retain their forms. However, we can reformulate the diffusion term:

### v2

@@ 6.2 Straight Trajectories
DDPM (${\mathcal{L}_{\mathrm{simple}}}$) [19] $1000$ $3.17$
DDPM (ELBO) [19] $1000$ $13.51$
Flow Matching [32] $142$ $6.35$ $122$ $5.02$ $138$ $14.14$
DDIM [56] $10$ $13.36$
@@
DPM Solver [37] $24$ $2.75$
Trajectory Curvature Minimization [31] $5$ $18.74$
@@
Multisample Flow Matching [43] $12$ $7.18$ $12$ $17.6$
@@
Figure 2: Trajectories of NFBM between two-dimensional data distributions, trained with an additional penalty to avoid obstacles.
6.3 Bridges
Table 3: FID values of dog $\rightarrow$ cat on AFHQ $64$.
Model FID $\downarrow$
DSBM [53] $14.16$
GSBM [33] $12.39$
NFBM $\mathbf{12.06}$
We consider the task of translating images of dogs into images of cats from the downsampled AFHQ dataset. Table 3 reports the FID scores on images sampled with $1000$ steps. Notably, NFBM outperforms the baselines, demonstrating the effectiveness of our framework for learning bridges.
Finally, we demonstrate that the framework flexibility also allows for learning dynamics with specific properties in the case of NFBM. To illustrate this, we adopted the strategy discussed in Section 5 and learn NFBM with an additional penalty to avoid obstacles (see details in Section F.2). Figure 2 visualizes the learned stochastic trajectories. It is evident that NFBM has efficiently learned to translate distributions while avoiding obstacles.
@@ 7 Related Work
Diffusion models, originally proposed by [55], have evolved significantly through subsequent developments [59, 19]. These advancements have resulted in remarkable generative quality for high-dimensional data distributions [12, 47]. Nevertheless, conventional diffusion models, typically relying on a linear Gaussian forward process, may not optimally fit some data distributions. To address this, alternative forward processes have been explored, such as combining blurring with Gaussian noise injection [45, 9, 22], diffusion in the wavelet spectrum [42], and forward processes based on the exponential family [40]. These models are limited by their fixed forward processes and may be seen as specific instances of NFDM.
@@
Some studies have focused on making the forward process learnable. Approaches include a learnable noise injection schedule [27, 38, 48] and learning data transformations like time-independent transformations based on VAEs [65, 46] and normalizing flows [25], or time-dependent transformations [17, 54, 39, 4]. These methods can also be considered special cases of NFDM with specific transformations $F_{\varphi}$ (eq. 7).
@@
To improve the sampling efficiency, in orthogonal line of works alternative sampling methods have been studied [61, 34, 37, 52]. Additionally, distillation techniques have been applied [49, 57, 73, 71] to enhance sampling speed, albeit at the cost of training additional models. Most of these approches are compatible with NFDM and may be combined for further gains.
Another line of works proposed technics for learning straighter generative trajectories. However, these approaches requires distillation [36, 35], solving unstable min-max problem [1], estimating an optimal coupling over minibatches [43, 63] of the entire dataset, which, for large datasets, may become uninformative, or can be viewed as specific instances [31, 51] of NFDM.
The connections between NFDM and some related works is discussed further in Appendix E.
8 Conclusion and Limitations
@@
In this paper we introduced NFDM, a novel simulation-free framework for improved diffusion modeling through a learnable forward processes. NFDM outperforms baseline diffusion models on standard benchmarks, showcasing its effectiveness. However, NFDM is not merely just another model, but rather it is a versatile framework that facilitates the predefined specification and learning of a forward process. Similarly, the penalty on the curvature of deterministic trajectories discussed in Section 5 represents just one specific example of the restrictions that NFDM can accommodate.
@@
Nonetheless, the advantages of NFDM come with certain trade-offs. Once the forward process is parameterized using a neural network, this leads to approximately 2.2 times longer optimization iteration of NFDM compared to conventional diffusion models. Additionally, as discussed in Section 3.1, NFDM doesn’t allow parameterisation of the forward process with functions, that doesen’t match certain conditions.
@@
Despite these challenges, we are optimistic about the potential of NFDM. NFDM opens up significant prospects for exploring new generative dynamics. We believe that with alternative parameterizations, modifications of the objective, and the integration of orthogonal approaches, NFDM has the potential to achieve even better results across a range of data modalities.
@@ Appendix A Derivations
A.1 Derivation of NFDM’s Objective
Our goal is to derive a variational bound on the negative log-likelihood of NFDM. For the convenience of the following derivations we will consider the forward process as a conditional reverse SDE (eq. 10) that starts from $\mathbf{z}_{1}\sim q_{\varphi}(\mathbf{z}_{1}|\mathbf{x})$ and flows backwards in time:
$d\mathbf{z}_{t}$ $=\tilde{f}^{B}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})dt+g_{\varphi}(t)d\bar{\mathbf{w}}.$ (17)
Similarly to a conditional SDE (eq. 9) flowing forwards in time, this SDE also corresponds to conditional marginal distribution $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$, which is implicitly defined by the invertible transformation $\mathbf{z}_{t}=F_{\varphi}(\varepsilon,t,\mathbf{x})$ (eq. 7), where $\varepsilon\sim q(\varepsilon)$.
For the reverse process we use the definition from Section 3.2 as reverse SDE (eq. 11) that starts from the prior distribution $p(\mathbf{z}_{1})$, flows backwards in time, and finally samples $\mathbf{x}\sim p(\mathbf{x}|\mathbf{z}_{0})$:
@@ A.1 Derivation of NFDM’s Objective
To derive the NFDM’s variational bound on the NLL we discretize processes, derive an objective in discrete case and then consider the limit of discretization to switch back to continuous time.
We discretize the conditional reverse SDE in eq. 17 and the reverse SDE in eq. 18, transitioning from continuous-time trajectories $\{\mathbf{z}(t)\}_{t\in[0,1]}$ to discrete-time trajectories $\bar{\mathbf{z}}_{0},\bar{\mathbf{z}}_{\frac{1}{T}},\dots,\bar{\mathbf{z}}_{1}$, where $T$ represents the number of discrete steps of size $\Delta t=\frac{1}{T}$.
The discrete conditional reverse process starts from distribution $\bar{q}_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})=q_{\varphi}(\bar{\mathbf{z}}_{1}|\mathbf{x})$ and follows conditional reverse Markov chain:
$\bar{q}_{\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t},\mathbf{x})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\tilde{f}^{B}_{\varphi}(\bar{\mathbf{z}}_{t},t,\mathbf{x}),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (19)
Similarly, the discrete generative process shares distributions $\bar{p}(\bar{\mathbf{z}}_{1})=p(\bar{\mathbf{z}}_{1})$ and $\bar{p}(\mathbf{x}|\bar{\mathbf{z}}_{0})=p(\mathbf{x}|\bar{\mathbf{z}}_{0})$ with the continuous one. However, instead of following the reverse SDE (eq. 18) it follows the reverse Markov chain:
$\bar{p}_{\theta,\varphi}(\bar{\mathbf{z}}_{t-\Delta t}|\bar{\mathbf{z}}_{t})=\mathcal{N}\Big{(}\bar{\mathbf{z}}_{t-\Delta t};\bar{\mathbf{z}}_{t}-\Delta t\hat{f}_{\theta,\varphi}(\bar{\mathbf{z}}_{t},t),\Delta tg^{2}_{\varphi}(t)I\Big{)}.$ (20)
@@
Now, considering these discrete processes and recognizing that they are Markovian conditional and unconditional processes moving backwards in time, we can leverage a key result from [55, 19]:
$\mathbb{E}_{q(\mathbf{x})}\left[-\log\bar{p}_{\theta,\varphi}(\mathbf{x})\right]\leq{\bar{\mathcal{L}}_{\mathrm{rec}}}+{\bar{\mathcal{L}}_{\mathrm{diff}}}+{\bar{\mathcal{L}}_{\mathrm{prior}}},\quad\textrm{where}$ (21)
¯${\bar{\mathcal{L}}_{\mathrm{rec}}}=\mathbb{E}_{\bar{q}_{\varphi}(\mathbf{x},\bar{\mathbf{z}}_{0})}\left[-\log p(\mathbf{x}|\bar{\mathbf{z}}_{0})\right],$ (22)
@@
Equation 21 establishes a variational upper bound on the negative log-likelihood of discretized reverse process. Taking the limits of this upper bound, the reconstruction term ${\mathcal{L}_{\mathrm{rec}}}={\bar{\mathcal{L}}_{\mathrm{rec}}}$ and the prior term ${\mathcal{L}_{\mathrm{prior}}}={\bar{\mathcal{L}}_{\mathrm{prior}}}$ retain their forms. However, we can rewrite the diffusion term: